Operating Systems and Systems Integration

Assignment 1: Redirection

Hand in one single page (can be handwritten, or you can send it by email if it is not a Microsoft format document).

Take some care with your answers.

1. Define each of the following terms:

(a) standard input

(b) standard output

(c) standard error

2. Explain what each of the following redirection operations are; briefly explain what they are used for, and what their purpose is:

(a) >

(b) 1>

(c) 2>

(d) >>

(e) 1>>
(f) 2>>

(g) >&

(h) <

(i) 0<

(j) |

Nick Urbanik <nicku@nicku.org>  ver. 1.1